Carnes Cosmetics Co.'s stock price is $47, and it recently paid a $1.50 dividend. This dividend is expected to grow by 30% for the next 3 years, then grow forever at a constant rate, g; and rs = 12%. At what constant rate is the stock expected to grow after Year 3? Do not round intermediate calculations. Round your answer to two decimal places.

  %
Answer #1

Solution: hiven. Do = 1.5 growth = 30% for next 34 Di: 1.5(1-3) = 1.95 D2 = 1.95 (1-3) - 2.535 D3 - 2-535(1-3) = 3.2955 D3 (9